The function f: R→R defined by f(x)= | x-1|+ | 3-x| is differentiable at x= 4.


True or false with full explanation

1
Expert's answer
2022-02-02T09:38:33-0500

The given function is "f(x)=|x-1|+|3-x|,\ x\in\R."


But, the derivative of a sum is the sum of derivatives so;

"\displaystyle
\frac{d}{dx} \left(\left|{x - 3}\right| + \left|{x - 1}\right|\right) = \frac{d}{dx} \left(\left|{x - 3}\right|\right) + \frac{d}{dx} \left(\left|{x - 1}\right|\right)"


The function "\displaystyle
\left|{x - 3}\right|" is the composition "\displaystyle
f{\left(g{\left(x \right)} \right)}" of two functions "\displaystyle
f{\left(u \right)} = \left|{u}\right|" and

"\displaystyle
g{\left(x \right)} = x - 3".

Also, The function "\displaystyle
\left|{x - 1}\right|" is the composition  "\displaystyle
h{\left(k{\left(x \right)} \right)}" of two functions "\displaystyle
h{\left(v \right)} = \left|{v}\right|" and

"\displaystyle
k{\left(x \right)} = x - 1"

Now, apply the chain rule which is "\displaystyle
\frac{d}{dx} \left(f{\left(g{\left(x \right)} \right)}\right) = \frac{d}{du} \left(f{\left(u \right)}\right) \frac{d}{dx} \left(g{\left(x \right)}\right)" and

"\displaystyle
\frac{d}{dx} \left(h{\left(k{\left(x \right)} \right)}\right) = \frac{d}{du} \left(h{\left(v \right)}\right) \frac{d}{dx} \left(k{\left(x \right)}\right)".

Where "\displaystyle
\frac{d}{du} \left(\left|{u}\right|\right) = \frac{u}{\left|{u}\right|}\text{ and }
\frac{d}{dv} \left(\left|{v}\right|\right) = \frac{v}{\left|{v}\right|}"

We have;

"\displaystyle
\frac{d}{dx} \left(\left|{x - 3}\right|\right) + \frac{d}{dx} \left(\left|{x - 1}\right|\right) = \frac{d}{du} \left(\left|{u}\right|\right) \frac{d}{dx} \left(x - 3\right) + \frac{d}{dv}(|v|)\frac{d}{dx} \left({x - 1}\right)"


"\displaystyle
=\frac{\left(u\right) \frac{d}{dx} \left(x - 3\right)}{\left|u\right|} + \frac{\left(v\right) \frac{d}{dx} \left(x - 1\right)}{\left|v\right|} \\"

"\displaystyle
=\frac{\left(x - 3\right) \frac{d}{dx} \left(x - 3\right)}{\left|{x - 3}\right|} + \frac{\left(x - 1\right) \frac{d}{dx} \left(x - 1\right)}{\left|{x - 1}\right|} \\"

"\displaystyle
 = \frac{\left(x - 3\right) }{\left|{x - 3}\right|} + \frac{x - 1}{\left|{x - 1}\right|}"

Thus,

"\displaystyle
\frac{d}{dx} \left(\left|{x - 3}\right| + \left|{x - 1}\right|\right) = \frac{x - 3}{\left|{x - 3}\right|} + \frac{x - 1}{\left|{x - 1}\right|}"

Hence, the given function is differentiable at "\displaystyle x=4" and

"\displaystyle
\frac{d}{dx} \left(\left|{x - 3}\right| + \left|{x - 1}\right|\right)|_{\left(x = 4\right)} = 2"